OpenStudy (anonymous):

If a storm window has an area of 400 square inches, what are the dimensions if the window is 12 inches higher than it is wide (w x h)?

OpenStudy (anonymous):

The windows being rectangular right?

OpenStudy (saifoo.khan):

x(x+12) = 400

OpenStudy (saifoo.khan):

x^2+12x -400 = 0

OpenStudy (saifoo.khan):

x = 14.88 x = -26.88

OpenStudy (saifoo.khan):

Area = b * h 400 = 14.88 * h h = 400/14.88 h = 26.88units